# Verifying Device Drivers with Pancake

### v1

@@ 2.5 Verification Approaches
Our aim is high-assurance machine code implementing device drivers. The verified compiler (Section 2.3) guarantees that our source code (written in Pancake) compiles to machine code that implements the source’s behaviour. This is half of the story: we must also be sure that the Pancake code we have compiled implements the device driver correctly (incorrect source $+$ verified compiler $=$ incorrect machine code).
The gold-standard (but expensive!) way to verify source code is to use interactive theorem-proving (ITP), as done by (for example) the seL4 [Klein et al., 2014] and CertiKOS [Gu et al., 2016] kernels, and the CompCert [Leroy, 2009], and CakeML [Kumar et al., 2014] (and thus Pancake) compilers. As Pancake is an imperative language, it would be natural to use an ITP-based implementation of Hoare logic as the foundation for the necessary proofs. We describe preliminary and future work on ITP-based approaches to source verification in Section 6.2.2 below.
However, as described in Section 3, our approach here is to use automated deductive verification, rather than an ITP, for productivity reasons. Using Viper [Müller et al., 2016a] (and its realisation of Hoare logic) increases the size of the trusted computing base (see Section 6.1 below), but can be expected to be less expensive in terms of developer time and expertise. The example of our case study bears this out.
Supporting this, Pancake syntax includes Viper annotations (visible in 1). These annotations resemble comments in the source code, and can appear at the top level and within function bodies. Top-level annotations are used, for instance, to specify function contracts, while in-function annotations can be used to specify loop invariants. The Pancake compiler treats annotations as null statements, and they disappear in the first phase of compilation. However, the Viper transpiler (see below) preserves the annotations so that verification can be carried out.
@@ 3 A Viper Front-end for Pancake
Our transpiler first uses the --explore option of the Pancake compiler to extract the abstract syntax tree (AST) of the input program, which includes Pancake-level (1) annotations of the kind shown in 1 as well as (2) hooks into a device model that we will describe in Section 3.3. It then translates that AST into encodings we have chosen for Pancake’s variables, values, memory locations, and annotated logical assertions over these, as expressed in Viper IL code.
The rest of this section will document and explain the most interesting of these encoding decisions, which regard our encoding of Pancake’s machine word type and memory in Viper. Our objective is to choose an encoding that maximises the performance of the resulting queries to Viper’s backend, while still being sound—that is, it does not produce a Viper query that is verifiable as true when the assertion from Pancake level that it is supposed to represent is actually false.
@@ 3.1 Machine-Word Encoding
As a language whose only primitive type is machine-size words (used for both values and pointers), Pancake’s word variables have bitvector semantics and overflows are well-defined behaviour, i.e., the variable is wrapped modulo the word size.
To preserve these semantics when encoding these as integers in Viper, which are signed and unbounded, one option is to treat arithmetic operations as modulo the word size. However, this approach significantly slows verification due to the mixing of arithmetic that is linear (such as addition, subtraction etc.) versus non-linear (such as modulo and bitvector operations) with respect to Viper’s unbounded signed integer space. Poor query performance due to mixing of different theories that require handling by different solver strategies is a well-known hazard for SMT solvers [Jovanovic and Barrett, 2013].
@@
Another concern is that bitvector operations are also non-linear with respect to Viper integers, causing similar performance concerns. Observing that almost all bitvector operations in our driver occur as part of bit masking, shifting etc. for accesses to device memory as part of a well defined device interface, we instead abstract this device interface, which we cover in Section 3.3. Additionally we precompute constant expressions and apply heuristics to rewrite common bitvector operations, e.g. x&255 is rewritten as x%256. Although the rewritten operations still use non-linear arithmetic, performance is improved compared to bitvector operations. This eliminates the vast majority of bitvector operations, with the remaining few in our Ethernet driver turning out to be reasonably performant.
@@ 3.2 Local Memory Encoding
As mentioned in Section 2, Pancake disallows pointers to stack variables, which simplifies modelling them in Viper, as for these we do not have have to manage any kind of access permission or to check for references to invalid memory.
For all other memory, the memory model of Pancake poses unique challenges due to its assembly-like nature, such as the lack of first-order support for arrays and reliance on pointer arithmetic for memory operations. Due to the lack of information about what different memory regions represent—be it an array, struct or other data structures—we must adopt the naivest modelling approach: with the exception of shared and device memory (described in the next section), we encode memory as an array of words. This approach results in non-idiomatic Viper code but captures Pancake’s native word-size treatment memory operations accurately.
The various memory regions a function needs access to, and with which permission, are added via annotations. These annotations are encoded as an iterated separating conjunction [Müller et al., 2016b], which can be verified efficiently in Viper.
@@ 3.3 Shared and Device Memory Encoding
As mentioned in Section 2.4, Pancake includes dedicated load/store operations for shared memory regions, whose behaviour in a sense resembles the effect of the volatile keyword in C which prevents the compiler from reordering or optimizing away those accesses. Thus, our driver uses these primitives for accesses to device memory, as well as accesses to memory shared between it and other OS components. Moreover, unlike the rest of memory, we cannot encode shared memory in Viper simply as an array of words, because we cannot rely on its contents not to change between accesses.
@@
This way, the driver developer does not necessarily have to specify the model of the device or neighbouring component before they implement their driver in Pancake, and furthermore they do not have to modify their Pancake driver’s shared memory accesses after specifying these models—they just add extra annotations to specify which Viper methods the transpiler should produce in place of loads/stores to given shared memory addresses. This approach ensures the separation of driver implementation and device specification, whilst also improving verification speeds.
@@ 4 Verified Ethernet Driver
We formally verify a single-core i.MX Ethernet driver for LionsOS [Heiser et al., 2025]. LionsOS uses a simple OS-side interface for device drivers using zero-copy shared-memory communication for network data, and lock-free, bounded, single-producer, single-consumer (SPSC) queues for meta data and control information. The driver synchronises with the rest of the OS via semaphores (implemented as seL4 Notifications).
@@
The target driver, implemented in Pancake, controls the MAC-NET 1 Gb/s Ethernet core common to NXP i.MX 8M Mini, Dual, QuadLite and Quad Applications Processors present in various Arm-based NXP system-on-chips (SoC), operating as a network interface card (NIC). The NIC uses DMA descriptor rings for passing the addresses of data buffers.
The driver incorporates formal specifications through annotations using Hoare logic and the Viper verification framework. This verification work establishes four classes of critical guarantees over the device driver, which we will examine in turn: (1) protocol compliance with NIC device interfaces (Section 4.1), (2) protocol compliance with neighbouring OS component interfaces (Section 4.2), (3) guaranteed data integrity across transfers (Section 4.3), and (4) memory safety through region separation and restricted access controls (Section 4.4).
4.1 Device Model and Properties
We model and enforce three kinds of properties with respect to the driver’s correct use of the device interfaces: (1) that device memory accesses are within the right address ranges, (2) that values written to the device obey the device’s requirements, and (3) that verification of the rest of the driver is robust for any values read from the device, as constrained only by the device model.
@@ 4.1 Device Model and Properties
Second, we verify that whenever the driver interacts with these device interfaces, it does so in the required way not to put the device into a bad state, as specified by its documentation and captured by our device model. To enforce this, we specify two kinds of requires clauses for device interface methods, as illustrated in 3 for a representative pair of examples, the store and load methods for EIR, a particular device register:
@@
Note that, while the presence of valid_device in the requires clauses of device interface methods requires the driver not to violate the validity of the device state, its presence also in the ensures clauses of all device methods specifies that we can assume the device itself will maintain that same validity invariant throughout all driver-device interactions.
Finally, we model the non-determinism of the values the driver could possibly obtain from the device by underspecifying its interface methods, forcing the verification of the calling context of the method (i.e. the Pancake code that invoked the shared memory operation) to account for a wide range of possible values constrained only by their ensures clauses. For example, load_EIR in 3 ensures that the return value is a valid unsigned 32-bit word value, as captured by bounded32(retval). Verification must subsequently succeed for any value returned by that interface that satisfies the ensures, i.e. any unsigned 32-bit word value.
@@ 4.2 OS Communication Protocols
Using much the same techniques just described in the previous section for specifying valid device interactions, we constrain the driver’s access to shared memory regions for network SPSC queues through annotations when interfacing with LionsOS components. We also model the shared SPSC queues non-deterministically to verify that the driver maintains protocol compliance—assuming the neighbouring OS component maintains it too—without assuming specific state values.
We verify that the driver adheres to the following network queue signaling protocols: When consuming data buffers from network queues, the driver requests wake-up signals from the OS when hardware rings have vacancy. When providing data buffers to the OS, the driver signals the OS’s semaphores if and only if (1) a signal was explicitly requested by the OS and (2) the queue state has changed by the driver [Heiser et al., 2025]. To prevent double signaling, we also ensure that the driver clears signal requests after notification. We also verify that queue states remain valid during all SPSC queue operations, for instance we check that no queue overflows or underflows occur at the driver side.
@@ 4.3 Data Integrity
We establish this integrity check on all pathways of data transfers in the driver. We also verify data transfer completeness by ensuring that within the driver, the number of SPSC queue operations align with the number of hardware descriptor ring state changes, so that there is no data loss in the driver.
@@ 4.4 Memory Access Control
In addition to the memory access constraints described in Section 4.1 and Section 4.2, we furthermore verify that only the parts of the driver responsible for packet transmission paths access any TX-related descriptor rings and SPSC queue state, and likewise that only its packet receipt paths access RX-related state.
We enforce this using Viper’s native permissions features to specify access controls, in this way providing formal verification of memory safety and region isolation. In effect, our driver verification applies a separation logic-like principle by partitioning the driver’s global memory into RX and TX regions (which reflects their use in LionsOS [Heiser et al., 2025]).
Figure 3: Driver verification workflow.
4.5 Verification Workflow
Figure 3 shows the resulting verification workflow. The verifier annotated the Pancake source of the driver and processes the resulting source with the transpiler, which produces the input for Viper. The verifier also supplies Viper specifications of the device interface, as well as the interface between the driver and the rest of LionsOS. These are then processed by Viper which either returns a result (proved or falsified) or times out.
@@ 5 Performance Evaluation
Our evaluation platform is an AVnet MaaXBoard with an NXP i.MX8MQ SoC, having four Arm Cortex A53 cores capable of a maximum of 1.5 GHz; we run our measurements at a fixed clock rate of 1 GHz. The board has 2 GiB of RAM and the on-chip 1 Gb/s NIC specified earlier (in Section 4).
@@
Figure 4 shows the result. The system has no problems handling the requested load: the received throughput equals the requested. The Pancake version of the driver uses slightly more CPU time to handle the load than the C version, the difference is about 10%.
Figure 5 puts this slight increase into context, by looking at latency (RTT) and the overall CPU use of the whole system. The difference in CPU usage of the drivers becomes unnoticeable. Similarly, the differences in latency is well within the variance of the results. We can summarise that the performance cost of the verified driver is in the noise.
@@ 6 Discussion
In Section 6.1, we discuss our TCB. Section 6.2 discusses alternative verification approaches we are pursuing that would mitigate the TCB. Finally, Section 6.3 discusses potential usability improvements to the verification front-end.
@@ 6.1 Threats to Validity
Moreover, while we plan to verify our transpiler from annotated Pancake to Viper (Section 6.2.1), this is not done yet. Until then, a mistake in the transpiler could produce a Viper query that the underlying SMT-based backend can prove true, even if the property as specified at the Pancake program point via Viper annotations is false. Beyond this, we also trust the Viper verification infrastructure and SMT solvers called by its backends to discharge the Hoare logic queries specified using its input language soundly, i.e. only when true.
@@
Finally, we trust the operating system kernel and its userland support libraries not to crash or malfunction. LionsOS runs on the seL4 OS microkernel, which is verified not to crash [Klein et al., 2014], and relies on the seL4 Microkit support library, whose main server loop has been verified not to exhibit undefined behaviour [Paturel et al., 2023]. However, functional properties of individual seL4 system calls relied on by Microkit have not yet been verified.

### v2

@@ 2.5 Verification Approaches
For the purpose of using Pancake to produce high-assurance device drivers, the verified compiler (Section 2.3) is only half of the story: we must also be sure that the Pancake code implements the device driver correctly (incorrect source $+$ verified compiler $=$ incorrect machine code).
The gold-standard (but potentially expensive) way to verify source code is to use interactive theorem-proving (ITP), as done by (for example) the seL4 [Klein et al., 2014] and CertiKOS [Gu et al., 2016] kernels, and the CompCert [Leroy, 2009], and CakeML [Kumar et al., 2014] (and thus Pancake) compilers. As Pancake is an imperative language, it would be natural to use an ITP-based implementation of Hoare logic. We describe preliminary and future work on ITP-based approaches to source verification in Section 6.2.
For productivity, given the commodity nature of devices, we use automated deductive verification, rather than an ITP. Using Viper [Müller et al., 2016b] (and its realisation of Hoare logic) increases the size of the trusted computing base (see Section 6.1), but can be expected to be less expensive in terms of developer time and expertise. Our case study bears this out.
To support this, Pancake syntax includes Viper annotations (visible in 1). They resemble comments in the source code, and can appear at the top level and within function bodies. Top-level annotations are used, for instance, to specify function contracts, while in-function annotations can be used for loop invariants. The Pancake compiler ignores annotations, but the Viper transpiler (Section 3) preserves the annotations so that verification can be carried out.
@@ 3 A Viper Front-end for Pancake
Our transpiler first uses a diagnostic feature of the Pancake compiler to extract the abstract syntax tree (AST) of the input program, which includes Pancake-level (1) annotations of the kind shown in 1 as well as (2) hooks into a device model that we will describe in Section 3.3. It then translates that AST into our chosen encodings for Pancake’s variables, values, memory locations, and annotated logical assertions over these, as expressed in Viper IL code.
The rest of this section documents and explains the most interesting of these encoding decisions, those concerning our encoding of Pancake’s machine word type, and our encoding of (shared and unshared) memory in Viper. Our objective an encoding that maximises the performance of the resulting queries to Viper’s backend, while still being sound—that is, it should never produce a Viper query that is verifiable as true, when the statement about Pancake code that it is supposed to encode is false.
@@ 3.1 Machine-Word Encoding
As a language whose only primitive type is machine-size words, Pancake’s word variables have bitvector semantics and overflows are well-defined behaviour, i.e., the variable is wrapped modulo the word size.
To preserve this semantics when encoding these as integers in Viper, which are signed and unbounded, one option is to treat arithmetic operations as modulo the word size. However, this approach significantly slows verification due to the mixing of arithmetic that is linear (such as addition, subtraction etc.) versus non-linear (such as modulo and bitvector operations) with respect to Viper’s unbounded signed integer space. Poor query performance due to mixing of different theories that require handling by different solver strategies is a well-known hazard for SMT solvers [Jovanovic and Barrett, 2013].
@@
Another concern is that bitvector operations are also non-linear with respect to Viper integers, causing similar performance concerns. Observing that almost all bitvector operations in our driver occur as part of bit masking, shifting etc., for accesses to device memory as part of a well defined device interface, we instead abstract this device interface, which we cover in Section 3.3. Additionally we precompute constant expressions and apply heuristics to rewrite common bitvector operations, e.g. x&255 is rewritten as x%256. Although the rewritten operations still use non-linear arithmetic, performance is improved compared to bitvector operations. This eliminates the vast majority of bitvector operations, with the remaining few in our Ethernet driver turning out to be reasonably performant as Viper queries.
@@ 3.2 Local Memory Encoding
As mentioned in Section 2, Pancake disallows pointers to stack variables, which simplifies modelling them in Viper: we do not have have to manage any kind of access permission or to check for references to invalid memory.
For all other memory, the memory model of Pancake poses unique challenges due to its assembly-like nature, such as the lack of first-order support for arrays and reliance on pointer arithmetic for memory operations. Due to the lack of information about what different memory regions represent—be they arrays, structs or other data structures—we must adopt the most naive modelling approach: with the exception of shared and device memory (described in the next section), we encode memory as an array of words. This approach results in non-idiomatic Viper code but captures Pancake’s native word-size treatment of memory operations accurately.
The various memory regions a function needs access to, and with which permission, are added via annotations encoded as an iterated separating conjunction [Müller et al., 2016a], which can be verified efficiently in Viper.
@@ 3.3 Shared and Device Memory Encoding
As mentioned in Section 2.4, Pancake includes dedicated load/store operations for shared memory regions, whose behaviour in a sense resembles volatile variables in C, whose accesses cannot be reordered or optimized away by the compiler. Thus, driver code uses these primitives for accesses to device memory, as well as accesses to memory shared between it and other OS components. Given this sharing, we cannot encode shared memory in Viper simply as an array of words, because we cannot rely on its contents not to change between accesses.
@@
This way, the driver developer is not required to specify the model of the device or neighbouring component before they implement their driver in Pancake. Furthermore, they do not have to modify their Pancake driver’s shared memory accesses after specifying these models—they instead add extra annotations to specify which Viper methods the transpiler should produce in place of loads/stores to given shared memory addresses. This approach ensures the separation of driver implementation and device specification, whilst also improving verification speeds.
@@ 4.1 Overview
We formally verify a single-core i.MX Ethernet driver for LionsOS [Heiser et al., 2025]. LionsOS uses a simple OS-side interface for device drivers with zero-copy shared-memory communication for network data. The interface employs lock-free, bounded, single-producer, single-consumer (SPSC) queues that contain meta data buffers for data addresses and lengths, as well as control information for signaling requests. The driver synchronises with the rest of the OS via semaphores (implemented as seL4 Notifications).
@@
The target driver, implemented in Pancake, controls the MAC-NET 1 Gb/s Ethernet core common to NXP i.MX 8M Mini, Dual, QuadLite and Quad Applications Processors present in various Arm-based NXP system-on-chips (SoC), operating as a network interface card (NIC). The NIC uses DMA descriptor rings for passing the addresses of data packets. This driver will be made available for LionsOS as a choice of network drivers alongside the C version. The driver incorporates formal specifications through annotations using Hoare logic and the Viper verification framework.
Figure 3: Driver verification workflow.
Component Line Count
C Driver 351
Pancake Driver 411
Pancake annotations 558
Viper Device model 391
Viper OS interfacing model 298
Table 1: Lines of code comparisons between the Ethernet driver in Pancake implementation and C implementation, and the lines of code of our verification artifacts.
Verification workflow
Figure 3 shows the resulting verification workflow. The programmer annotates the Pancake source of the driver and processes the resulting source with the transpiler, which produces the Viper input. The programmer also supplies Viper specifications of the device interface as well as the interface between the driver and the rest of LionsOS. These are then processed by Viper, which returns a result (proved or falsified) or times out. Table 1 shows the total size of our verification artifacts.
Device model
We model the hardware NIC device and its non-deterministic state in Viper specifications. The model is then connected to Pancake driver source code after transpilation, using the top-level Pancake annotations as described in Section 3.3.
The device state consists of RX (receive) and TX (transmit) hardware descriptor rings, with each ring decomposed into three integer sequences representing data addresses, lengths, and bitfields. These integers represent machine words and are converted to bitvectors when bit-level operations are required. As discussed above, by isolating the descriptor components, such a decomposition simplifies verification and reduces SMT solver complexity, particularly when using non-linear arithmetic and bitvector operations. Each shared memory operation corresponds to a Viper method in the device model, with memory loads returning non-deterministic values within valid ranges, and memory stores satisfying both address and value requirements.
We model the non-determinism of the values the driver could possibly obtain from the device by underspecifying its interface methods, forcing the verification of the calling context of the method (i.e. the Pancake code that invoked the shared memory operation) to account for a wide range of possible values constrained only by their ensures clauses. For example, load_EIR in 3 ensures that the return value is a valid unsigned 32-bit word value, as captured by bounded32(retval). Verification must subsequently succeed for any value returned by that interface that satisfies the ensures, i.e., any unsigned 32-bit word value.
What is verified
This verification work establishes the following four classes of critical guarantees:
  1. 1.
    Compliance with NIC device interfaces (Section 4.2): we verify that:
    • device memory accesses are within the right address ranges; and
    • values written to the device obey the device’s requirements.
  2. 2.
    Compliance with neighbouring OS component interfaces (Section 4.3): we verify that the driver adheres to the network queue signalling protocols, specifically:
    • for consuming meta data buffers from network queues, the driver requests the OS wake-up signals OS if hardware rings have vacancy,
    • for providing meta data buffers to the OS, the driver notifies the OS’s semaphores if and only if (i) a signal was explicitly requested by the OS and (ii) the queue state has been changed by the driver; and
    • the driver clears signal requests after notifying the OS (to prevent double signalling)
  3. 3.
    Guaranteed data integrity across transfers (Section 4.4): we verify that there is no data loss in the driver by enforcing integrity check on packet meta data (addresses and lengths) on all pathways of meta data transfers in the driver (note that, with LionsOS [Heiser et al., 2025], the driver never needs to access or map the packet data itself, and hence never violates the integrity or confidentiality of the packet data); and
  4. 4.
    Memory safety through region isolation and restricted access controls (Section 4.5): we verify that access to TX-related states and RX-related states are correctly restricted to the parts of the driver responsible for transmitting and receiving, respectively.
4.2 Device Interface Compliance
@@
Second, we verify that, whenever the driver interacts with these device interfaces, it does so in the required way not to put the device into a bad state, as specified by its documentation and captured by our device model. To enforce this, we specify two kinds of requires clauses for device interface methods, as illustrated in 3 for a representative pair of examples, the store and load methods for EIR, a particular device register:
@@
Note that, while the presence of valid_device in the requires clauses of device interface methods requires the driver not to violate the validity of the device state, its presence also in the ensures clauses of all device methods specifies that we can assume the device itself will maintain that same validity invariant through all driver-device interactions.
@@ 4.3 OS Communication Protocols
Using much the same approach just described in the previous section for specifying valid device interactions, we constrain the driver’s access to shared memory regions for network SPSC queues through annotations when interfacing with LionsOS components. We also model the shared SPSC queues non-deterministically to verify the driver maintains protocol compliance—assuming the neighbouring OS component maintains it too—without assuming specific state values.
@@ 4.4 Data Integrity
We establish this integrity check on all pathways of data transfers in the driver. We also verify data transfer completeness by ensuring that within the driver, the number of SPSC queue operations align with the number of hardware descriptor ring state changes, so there is no data loss in the driver.
@@ 4.5 Memory Access Control
In addition to the memory access constraints described in Sections 4.2 and 4.3, we also verify that only the parts of the driver responsible for packet transmission access any TX-related descriptor rings and SPSC queue state, and likewise that only its packet receipt paths access RX-related state.
We enforce this using Viper’s native permissions to specify access controls, in this way providing formal verification of memory safety and region isolation. In effect, our driver verification applies a separation logic-like principle by partitioning the driver’s global memory into RX and TX regions (which reflects their use in LionsOS [Heiser et al., 2025]).
@@ 5 Performance Evaluation
Our evaluation platform is an AVnet MaaXBoard with an NXP i.MX8MQ SoC, having four Arm Cortex A53 cores capable of a maximum of 1.5 GHz. We run our measurements at a fixed clock rate of 1 GHz to avoid the need for thermal management. [^2] Note that running at a low clock rate will magnify software overheads, and as such presents a pessimistic performance. The board has 2 GiB of RAM and the on-chip 1 Gb/s NIC specified earlier (in Section 4).
@@
Figure 4 shows the result. Like the C driver, the Pancake driver has no problems handling the requested load: the received throughput equals the requested. The Pancake version of the driver uses slightly more CPU time to handle the load than the C version, the difference is about 10%. Viewed in the context of LionsOS using less than half the CPU of Linux to handle a particular network load, the slightly increased overhead of the Pancake driver is in the noise.
Figure 5 puts this slight increase into context, by looking at latency (RTT) and the overall CPU use of the whole system. The difference in CPU usage of the drivers becomes unnoticeable. Similarly, the differences in latency is well within the variance of the results. This is in the context of LionsOS using less than half the CPU of Linux to handle a particular network load [Heiser et al., 2025]. We can summarise that the performance cost of the verified driver is in the noise.
@@ 6 Discussion
In Section 6.1, we discuss our TCB. Section 6.2 discusses alternative verification approaches we are pursuing that would reduce the TCB further. Finally, Section 6.3 discusses potential usability improvements to the verification front-end.
@@ 6.1 Threats to Validity
Moreover, while we plan to verify our transpiler from annotated Pancake to Viper (Section 6.2), this is not done yet. Until then, a mistake in the transpiler could produce a Viper query that the underlying SMT-based backend can prove true, even if the property as specified at the Pancake program point via Viper annotations is false. Beyond this, we also trust the Viper verification infrastructure and SMT solvers called by its backends to discharge the Hoare logic queries specified using its input language soundly, i.e., only when true.
@@
Finally, we trust the operating system kernel and its user-level support libraries not to crash or malfunction. LionsOS runs on the seL4 OS microkernel, which is verified not to crash [Klein et al., 2014], and relies on the seL4 Microkit support library, whose main server loop has been verified not to exhibit undefined behaviour [Paturel et al., 2023]. However, functional properties of individual seL4 system calls relied on by Microkit have not yet been verified.